OpenStudy (anonymous):

With 95% confidence , between which values do you think that the proportion of the adult population (35 years or older) of XXX City that favor lowering the drinking age to 18 is? A. 0.6  0.0635 B. 0.7  0.0635 C. 0.6  0.0679 D. 0.7  0.0679 E. 200  0.0679

OpenStudy (anonymous):

All the questions refer to the following story Please read the story carefully A certain organization wants to know the opinion of the adult residents, 35 years of age or older, of the 'XXX' City with a total population of 55,000 (similar size of Johnson City) about lowering the drinking age to 18. Specifically they want to know what proportion of those adults favor lowering the drinking age to 18.They organize a survey and select a simple random sample of 200 adult residents (35 years of age or older) and ask them "Do you think that the drinking age should be reduced to 18?" . In the sample, 140 people answered "yes".

OpenStudy (judygreeneyes):

do you still need an answer or is it too late?

OpenStudy (judygreeneyes):

B. 0.7 +/- 0.0635 is the answer. The confidence interval is \[p \pm 1.96 \sqrt{\frac{ p(1-p) }{ n }}\] p = 140/200 = 0.7 (the proportion in favor of lowering the drinking age) n = the sample size, 200 1.96 is the critical value for a 95% confidence interval, from the z-table (standard normal distribution)